What is the radius (in cm) of a cylindrical silicon block (density = 2.03 g/cm3) with the height of 3.20 cm to react with 76.0 g of sodium hydroxide by the reaction below? 

Si(s) + NaOH(aq) + H2O(l) → Na2SiO3(aq) + H2(g)


1
Expert's answer
2021-10-27T06:41:01-0400

"Si +4NaOH+2H_2O\to 2NaSiO_3+4H_2"

moles of NaOH="\frac{76}{23+16+1}=1.9"

mole ratio Si : NaOH=2:4=1:2

moles of "Si=1.9\/2=0.95"

mass of Si "=0.95\times 28=26.6g"

volume of silicon "=mass\/density=26.6\/2.03=13.1cm^3"

volume of cylindrical silicon "=\pi \times r^2 \times h=13.1"

"r^2=\frac{13.1}{\pi \times 3.2}=1.303"

"r=\sqrt{1.303} =1.14cm"